# NegotiationToM: A Benchmark for Stress-testing Machine Theory of Mind on Negotiation Surrounding

### v2

@@ 1 Introduction
The negotiation dialogues contain complicated and diverse aspects of a realistic negotiation, such as rapport building, discussing preferences, exchanging offers, emotional expression, and persuasion with personal and logical arguments Chawla et al. (2021b). In a realistic negotiation, humans innately infer the mental states of the other party and proceed with their subsequent actions based on their own beliefs and desires. For example, in Figure 1, two agents negotiate for food, water, and firewood packages for their upcoming trip. Initially, agent 1 lacks any information pertaining to the preference order of agent 2. Thus, agent 1’s belief is “no information on the agent 2 preference item”, and agent 1 intends to elicit the item preference order from agent 2 to guide further action based on their own belief. Furthermore, belief is commonly employed to denote an individual’s cognitive stance or acceptance of something as true or holding it to be the case Turiel (2008). This belief may undergo changes during the negotiation as perceiving more available information behind the conversation. In the fourth round of conversation is depicted in Figure 1, agent 1’s dynamic belief changed from “Agent 2’ high preference is food, medium preference is firewood, and the low one is water” to “Agent 2’s high preference is firewood, medium preference is food, and the low one is water.” Therefore, negotiation serves as an ideal scenario to assess the theory of mind ability of language learning models in the real world due to its complexity and the linguistic diversity inherent in negotiation conversations.
In this work, we introduce NegotiationToM, a natural conversational benchmark for stress-testing machine ToM in real-world negotiation surroundings involving multi-dimensional mental states (i.e., desires, beliefs, and intentions), inspired by the Belief-Desire-Intention (BDI) agent model proposed by Bratman (1987). The goal of NegotiationToM is to effectively measure how well large language models (LLMs) can track the mental states of negotiation participants in conversations and evaluate LLMs’ capability for a coherent understanding of others’ mental states in the conversation context where there are gradually increasing rounds of utterance (i.e., increase available information). We hope our benchmark and experimental results in the real-world scenario complement the prior theoretical works, which yield important insights into the intensive debate around ToM Whang (2023) in LLMs. We will release our benchmark and LLMs’ response to spark further discussions on evaluating the ToM capabilities of LLMs. Our contributions are summarized as follows:
@@ 3 NegotiationToM
Theory of Mind (ToM) describes the ability as humans have to ascribe and infer the mental states of others, and to predict which likely actions they are going to take Apperly (2010). Therefore, it is critical to acquire negotiation strategies based on one’s own desire and temporary belief built upon the information presented in conversation. Nevertheless, understanding the Theory of Mind (ToM) inherent in a negotiation dialogue is challenging due to its intricate linguistic features and complex reasoning attributes. Therefore, there are some considerations that have to be taken into account when constructing the NegotiationToM.
@@ 3.1 Design Considerations for NegotiationToM
There are several essential design considerations we go through when constructing the NegotiationToM. (1) the scenario of the dataset should be grounded in a human-to-human real-world negotiation (e.g., real-world camping scenario). (2) the dataset should be a natural conversational dataset instead of generated from a synthetic template to avoid reporting bias Gordon and Durme (2013) and shortcuts Aru et al. (2023). (3) the dataset should be equipped with abundant and diverse linguistic features and semantic context (e.g., negotiation argument) instead of bargaining on the numerical value or meaningless counter-offer Lewis et al. (2017); He et al. (2018).
@@ 3.2 CaSiNo
Inspired by several considerations above, CaSiNo Chawla et al. (2021b) was employed as the source data and modified to construct NegotiationToM. CaSiNo is a bilateral human-to-human natural conversational dataset that covers rich linguistic features and many realistic aspects of negotiations, such as small talk, preference elicitation, emotional expression, and convincing strategies based on individual desire. In this dataset, the participants take the role of campsite neighbors and negotiate for food, water, and firewood packages for their upcoming trip. For each conversation, participants discuss individual needs by making various convincing arguments from their camping experiences, such as Personal Care, Recreational, Group Needs, or Emergency requirements. One example of Group Needs is "I need more firewood due to having several people join on the trip and needing a bigger fire overall." We illustrate some of these arguments in Table 21 in Appendix A.2. Therefore, crafting our benchmark from the CaSiNo offers a range of scenarios based on how to align the preferences of the two parties to reveal more interesting behaviors.
@@ 3.3 Theory of Mind in NegotiationToM
Intention is a mental state formed through rational planning (i.e., negotiation strategy in a negotiation scenario) toward a goal based on the desires and beliefs of the agent. Intentions have been extensively explored in psychology tests, e.g., action prediction Malle and Knobe (2001) and intention attribution to abstract figures Castelli (2006). Normally, a negotiation strategy is highly associated with corresponding concrete intentions Belmondo and Sargis-Roussel (2015). Thus, in NegotiationToM, we collect the annotated negotiation strategies from the CaSiNo dataset and map the intentions according to the definition of various strategies. The mapping table is shown in Table 2, and the strategy definition and examples are illustrated in Table 19 and 20 in Appendix A.2. Within our framework, as both Self-Need and Other-Need are associated with "Intents to describe a need for an item" intention, we combine these two strategies into one intention class.
@@ 3.4 Annotation & Statistics
NegotiationToM is annotated based on a multi-turn negotiation dialogue corpus, the CaSiNo Chawla et al. (2021a) dataset. Each instance in CaSiNo is an -round alternating dialogue between two participants, and 222When the dialogue ends with user , is an empty utterance.. They take on the roles of campsite neighbors and negotiate for food, water, and firewood packages for their upcoming trip. We adopt the subset with strategy annotations and undertake the annotation on the desire and belief states behind each utterance..
@@
The intention state of both participants has already been introduced and mapped from the strategy annotations in CaSiNo. We conduct an expert annotation to annotate the beliefs and desires of the two participants in each dialogue (i.e., the perceived preference ranking among food, water, and firewood). We recruited five workers who were graduate students in English-speaking universities to conduct the annotation. For each dialogue , let be the truncated dialogue until round : . Then, we ask the workers to annotate the perceived preference ranking for both participants and for truncated dialogue () given all rounds of historical dialogue.
To ensure the annotation quality, we evaluate the workers on the first 100 rounds of conversations and explain to them in detail about their typical errors. More details of the annotation process are in Appendix A.1. Although annotating NegotiationToM requires understanding complex dialogues in CaSiNo, we observed high inter-annotator agreement. The Fleiss’s is 79.03% Fleiss (1971) for NegotiationToM benchmark, the breakdown computation of are shown in Table 3.
@@ 4.1 Baseline Models
In this work, we test six recent instruction-tuned large language models: GPT-4 OpenAI (2023), ChatGPT (OpenAI, 2022), Claude-v1.3333https://www.anthropic.com/news/introducing-claude Anthropic (2023a), Claude-v2.1444https://www.anthropic.com/news/claude-2-1 Anthropic (2023b), Llama-2 Chat 13B Touvron et al. (2023), and Llama-2 Chat 70B Touvron et al. (2023). Descriptions for each model are in Appendix B.1. By following the common practices in the theory of mind field Kim et al. (2023); Gandhi et al. (2023); Shapira et al. (2023b), we test these models with two types of prompts: (1) one is zero-shot prompting and we utilize the prompt template in Robinson and Wingate (2023) to formulate the task as a multiple choice question answering problem as a baseline. (2) another one is the Chain-of-Thought (CoT) prompting method by following Kojima et al. (2022) and using the prompt “let’s think step by step.” Apart from these two settings, we also assess the LLMs’ performance by using the few-shot setting to validate whether LLMs can improve their performance with input-output exemplars. We append four tailored exemplars for desire and belief states and seven designed exemplars for intentions, covering all the input-output exemplars’ labels. More configuration details can be found in Appendix B.2, and the prompt template refers to Appendix B.4. To measure the specific performance gap between humans and the state-of-the-art machine on the NegotiationToM, we employ three graduate students in computer science to complete the human evaluation task. More details of human evaluation are shown in Appendix C.2.
@@ 5.4 The Effect of Question Format
With the performance of LLMs varying significantly due to the sensitivity of prompt templates Webson and Pavlick (2022), we assessed the performance of two state-of-the-art models, ChatGPT and GPT-4, to study the effect of various question formats on their performance. In this experiment, we adopt three types of question formats, including ranking format, individual format, and combined format for desire and belief mental state. The combined format is the baseline prompt template adopted in our main experiment, which combines all three questions regarding the preference order into a single question and asks the LLMs to answer it simultaneously. The ranking format indicates collecting high, medium, and low preference items as one ranking answer. The individual format splits the high, medium, and low preference questions into three questions and feeds them to LLMs individually. The combined, ranking, individual, question format prompt template is shown in Tables 8, 13, 14, and 15 in Appendix B.4.
@@ 5.5 CaSiNo Negotiation Strategy Prediction
To validate the significance of our annotated desire and belief states, we append the information from these two states into the prompt template and assess whether it enhances the model performance on the negotiation strategy prediction task from the CaSiNo dataset. The baseline prompt template and the prompt template with belief and desire are shown in Tables 16 and 17 in Appendix B.4. The micro F1 score and macro F1 were employed as this task metric, and the result is reported in Table 6. All models incorporating the information from these two mental states received a significant improvement over the baselines. Specifically, by integrating the signals from desire and belief, the Claude-v2.1 model obtained a 7.44% gain in the micro F1 score and a 5.54% gain in the macro F1 score. It demonstrates that the effectiveness of our annotated theory of mind states (i.e., desire and belief) helps LLMs to infer the negotiation strategy behind each utterance. For example, with the understanding that agent 1’s preference order of items is Not Given, Not Given, and Not Given, and agent 2’s preference order of items is Firewood, Not Given, and Not Given. Agent 2 may take the elicit-preference strategy to elicit the preference order of agent 1 for further negotiation.
@@ 6 Conclusion
This work introduces NegotiationToM, a new benchmark designed to stress-test machine ToM in real-world negotiation surrounding covered multi-dimensional mental states. We perform extensive in-depth experiments to evaluate LLMs’ capability on the NegotiationToM benchmark.
@@
In this work, we conformed to recognized privacy practices and rigorously followed the data usage policy. We declare that all authors of this paper acknowledge the ACM Code of Ethics and honor the code of conduct. This paper introduces a benchmark for stress-testing machine theory of mind of large language model on the negotiation surrounding. Our benchmark is modified from the CaSiNo Chawla et al. (2021b), an English-based negotiation dataset. They conducted a data post-processing step for filtering inappropriate language use (e.g., English swear words) dialogues although this situation rarely occurred in the negotiation process. Therefore, we can foresee no immediate social consequences or ethical issues as we do not introduce social/ethical bias into the model or amplify any bias from the data. Moreover, the license CaSiNo CC-BY-4.0 license allows us to modify the data for research, and this fulfills their intended use.
@@ Appendix A Appendix for NegotiationToM
A.1 NegotiationToM Annotation
@@
A.2 Details for NegotiationToM Intentions
@@ B.4 Appendix for Prompt Template
In this section, we introduce all prompt template and these template are presented in Tables 8, 9, 10, 11, 12, 13, 14, 15, 16, and 17.
@@
Table 8 introduces the baseline prompt template, a straightforward Q&A session without requiring detailed and complex explanations. Table 9 is the Chain-of-Thought prompt template that facilitates a detailed, step-by-step thought process in the LLM, thoroughly considering each problem aspect before arriving at a solution. Tables 10, 11, and 12 are presented the few-shot setting prompt template. Each prompt template appends some exemplars for LLMs to learn the new tasks.
@@
The Ranking Question Format in Table 13 asks the LLM to prioritize or rank items, which is useful for tasks needing decision-making based on preference or importance. Table 14 and 15, the individual question format prompt templates, break down desire and belief questions into separate and focused questions.
@@
Tables 16 and 17 present the prompt template for the strategy prediction task; the former provides a baseline format for predicting strategies from negotiation conversations, while the latter enriches this task by incorporating the desires and beliefs of the agents involved, offering a deeper contextual understanding and offer signals from desire and belief states for strategy classification.

### v3

@@ 1 Introduction
The negotiation dialogues contain complicated and diverse aspects of a realistic negotiation, such as rapport building, discussing preferences, exchanging offers, emotional expression, and persuasion with personal and logical arguments Chawla et al. (2021). In a realistic negotiation, humans innately infer the mental states of the other party and proceed with their subsequent actions based on their own beliefs and desires. For example, in Figure 1, two agents negotiate for food, water, and firewood packages for their upcoming trip. Initially, agent 1 lacks any information pertaining to the preference order of agent 2. Thus, agent 1’s belief is “no information on the agent 2 preference item”, and agent 1 intends to elicit the item preference order from agent 2 to guide further action based on their own belief. Furthermore, belief is commonly employed to denote an individual’s cognitive stance or acceptance of something as true or holding it to be the case Turiel (2008). This belief may undergo changes during the negotiation as perceiving more available information behind the conversation. In the fourth round of conversation depicted in Figure 1, agent 1’s dynamic belief changed from “Agent 2’ high preference is food, medium preference is firewood, and the low one is water” to “Agent 2’s high preference is firewood, medium preference is food, and the low one is water.” Therefore, negotiation serves as an ideal scenario to assess the theory of mind ability of language learning models in the real world due to its complexity and the linguistic diversity inherent in negotiation conversations.
In this work, we introduce NegotiationToM, a natural conversational benchmark for stress-testing machine ToM in real-world negotiation surroundings involving multi-dimensional mental states (i.e., desires, beliefs, and intentions), inspired by the Belief-Desire-Intention (BDI) agent model proposed by Bratman (1987). The goal of NegotiationToM is to effectively measure how well large language models (LLMs) can track the mental states of negotiation participants in conversations and evaluate LLMs’ capability for a coherent understanding of others’ mental states in the conversation context where there are gradually increasing rounds of utterance (i.e., increase available information). We hope our benchmark and experimental results in the real-world scenario complement the prior theoretical works, which yield important insights into the intensive debate around ToM Whang (2023) in LLMs. Our contributions are summarized as follows:
@@ 3 NegotiationToM
Theory of Mind (ToM) describes the ability as humans have to ascribe and infer the mental states of others, and to predict which likely actions they are going to take Apperly (2010). Therefore, it is critical to acquire negotiation strategies based on one’s own desire and temporary belief built upon the information presented in conversation. Nevertheless, understanding the Theory of Mind (ToM) inherent in a negotiation dialogue is challenging due to its intricate linguistic features and complex reasoning attributes. Therefore, some considerations have to be taken into account when constructing the NegotiationToM.
@@ 3.1 Design Considerations for NegotiationToM
There are several essential design considerations we go through when constructing the NegotiationToM. (1) the scenario of the dataset should be grounded in a human-to-human real-world negotiation (e.g., real-world camping scenario). (2) the dataset should be a natural conversational dataset instead of generated from a synthetic template to avoid reporting bias Gordon and Durme (2013) and shortcuts Aru et al. (2023). (3) the dataset should be equipped with abundant and diverse linguistic features and semantic context (e.g., negotiation argument) instead of bargaining on the numerical value or meaningless counter-offer Lewis et al. (2017); He et al. (2018). (4) the dataset should be ensured to mitigate the risk of potential contamination.
@@ 3.2 CaSiNo
Inspired by several considerations above, CaSiNo Chawla et al. (2021) was employed as the source data and modified to construct NegotiationToM. CaSiNo is a bilateral human-to-human natural conversational dataset that covers rich linguistic features and many realistic aspects of negotiations, such as small talk, preference elicitation, emotional expression, and convincing strategies based on individual desire. In this dataset, the participants take the role of campsite neighbors and negotiate for food, water, and firewood packages for their upcoming trip. For each conversation, participants discuss individual needs by making various convincing arguments from their camping experiences, such as Personal Care, Recreational, Group Needs, or Emergency Requirements. One example of Group Needs is "I need more firewood due to having several people join on the trip and needing a bigger fire overall." We illustrate some of these arguments in Table 22 in Appendix A.3. Therefore, crafting our benchmark from the CaSiNo offers a range of scenarios based on how to align the preferences of the two parties to reveal more interesting behavior. Furthermore, we present the verification method employed to alleviate the risk of potential contamination within the CaSiNo dataset and demonstrate that this dataset is unlikely to encounter the contamination issue, as detailed in Appendix A.1.
@@ 3.3 Theory of Mind in NegotiationToM
Intention is a mental state formed through rational planning (i.e., negotiation strategy in a negotiation scenario) toward a goal based on the desires and beliefs of the agent. Intentions have been extensively explored in psychology tests, e.g., action prediction Malle and Knobe (2001) and intention attribution to abstract figures Castelli (2006). Normally, a negotiation strategy is highly associated with corresponding concrete intentions Belmondo and Sargis-Roussel (2015). Thus, in NegotiationToM, we collect the annotated negotiation strategies from the CaSiNo dataset and map the intentions according to the definition of various strategies. The mapping table is shown in Table 2, and the strategy definition and examples are illustrated in Table 20 and 21 in Appendix A.3. Within our framework, as both Self-Need and Other-Need are associated with "Intents to describe a need for an item" intention, we combine these two strategies into one intention class.
@@ 3.4 Annotation & Statistics
NegotiationToM is annotated based on a multi-turn negotiation dialogue corpus, the CaSiNo Chawla et al. (2021) dataset. Each instance in CaSiNo is an -round alternating dialogue between two participants, and 222When the dialogue ends with user , is an empty utterance.. They take on the roles of campsite neighbors and negotiate for food, water, and firewood packages for their upcoming trip. We adopt the subset with strategy annotations and undertake the annotation on the desire and belief states behind each utterance.
@@
The intention state of both participants has already been introduced and mapped from the strategy annotations in CaSiNo. We conduct an expert annotation to annotate the beliefs and desires of the two participants in each dialogue (i.e., the perceived preference ranking among food, water, and firewood). We recruited five workers who were graduate students in English-speaking universities to conduct the annotation. For each dialogue , let be the truncated dialogue until round : . Then, we ask the workers to annotate the perceived preference ranking for both participants and for truncated dialogue () given all rounds of historical dialogue. To ensure the annotation quality, we evaluate the workers during the first 100 rounds of conversations and explain their typical errors to them in detail. More details of the annotation process are in Appendix A.2. Although annotating NegotiationToM requires understanding complex dialogues in CaSiNo, we observed high inter-annotator agreement. The Fleiss’s is 79.03% Fleiss (1971) for NegotiationToM benchmark, the breakdown computation of are shown in Table 3.
@@ 4.1 Baseline Models
In this work, we test six recent instruction-tuned large language models: GPT-4 OpenAI (2023), ChatGPT (OpenAI, 2022), Claude-v1.3333https://www.anthropic.com/news/introducing-claude Anthropic (2023a), Claude-v2.1444https://www.anthropic.com/news/claude-2-1 Anthropic (2023b), Llama-2 Chat 13B Touvron et al. (2023), and Llama-2 Chat 70B Touvron et al. (2023). Descriptions for each model are in Appendix B.1. By following the common practices in the theory of mind field Kim et al. (2023); Gandhi et al. (2023); Shapira et al. (2023b), we test these models with two types of prompts: (1) one is zero-shot prompting and we utilize the prompt template in Robinson and Wingate (2023) to formulate the task as a multiple choice question answering problem as a baseline. (2) another one is the Chain-of-Thought (CoT) prompting method by following Kojima et al. (2022) and using the prompt “let’s think step by step.” Apart from these two settings, we also assess the LLMs’ performance by using the few-shot setting to validate whether LLMs can improve their performance with input-output exemplars. We concatenate four tailored exemplars for desire and belief states and seven designed exemplars for intentions, covering all the input-output exemplars’ labels. More configuration details can be found in Appendix B.2, and the prompt template refers to Appendix B.4. To measure the specific performance gap between humans and the state-of-the-art machine on the NegotiationToM, we employ three graduate students in computer science to complete the human evaluation task. More details of human evaluation are shown in Appendix C.2.
@@ 5.4 The Effect of Question Format
With the performance of LLMs varying significantly due to the sensitivity of prompt templates Webson and Pavlick (2022), we assessed the performance of two state-of-the-art models, ChatGPT and GPT-4, to study the effect of various question formats on their performance. In this experiment, we adopt three types of question formats, including ranking format, individual format, and combined format for desire and belief mental state. The combined format is the baseline prompt template adopted in our main experiment, which combines all three questions regarding the preference order into a single question and asks the LLMs to answer it simultaneously. The ranking format indicates collecting high, medium, and low preference items as one ranking answer. The individual format splits the high, medium, and low preference questions into three questions and feeds them to LLMs individually. The combined, ranking, individual, question format prompt template is shown in Tables 9, 14, 15, and 16 in Appendix B.4.
@@ 5.5 CaSiNo Negotiation Strategy Prediction
To validate the significance of our annotated desire and belief states, we append the information from these two states into the prompt template and assess whether it enhances the model performance on the negotiation strategy prediction task from the CaSiNo dataset. The baseline prompt template and the prompt template with belief and desire are shown in Tables 17 and 18 in Appendix B.4. The micro F1 score and macro F1 were employed as this task metric, and the result is reported in Table 6. All models incorporating the information from these two mental states received a significant improvement over the baselines. Specifically, by integrating the signals from desire and belief, the Claude-v2.1 model obtained a 7.44% gain in the micro F1 score and a 5.54% gain in the macro F1 score. It demonstrates that the effectiveness of our annotated theory of mind states (i.e., desire and belief) helps LLMs to infer the negotiation strategy behind each utterance. For example, with the understanding that agent 1’s preference order of items is Not Given, Not Given, and Not Given, and agent 2’s preference order of items is Firewood, Not Given, and Not Given. Agent 2 may take the elicit-preference strategy to elicit the preference order of agent 1 for further negotiation.
@@ 6 Conclusion
This work introduces NegotiationToM, a new benchmark designed to stress-test machine ToM in real-world negotiation surrounding covered multi-dimensional mental states. We performed comprehensive and detailed experiments to evaluate LLMs’ capability on the NegotiationToM benchmark and discovered that LLMs exhibit inferior performance compared to humans in the NegotiationToM task.
@@
In this work, we conformed to recognized privacy practices and rigorously followed the data usage policy. We declare that all authors of this paper acknowledge the ACM Code of Ethics and honor the code of conduct. This paper introduces a benchmark for stress-testing machine theory of mind of large language model on the negotiation surrounding. Our benchmark is modified from the CaSiNo Chawla et al. (2021), an English-based negotiation dataset. They conducted a data post-processing step for filtering inappropriate language use (e.g., English swear words) dialogues although this situation rarely occurred in the negotiation process. Therefore, we can foresee no immediate social consequences or ethical issues as we do not introduce social/ethical bias into the model or amplify any bias from the data. Moreover, the license CaSiNo CC-BY-4.0 license allows us to modify the data for research, and this fulfills their intended use.
Acknowledgements
The authors of this paper were supported by the NSFC Fund (U20B2053) from the NSFC of China, the RIF (R6020-19 and R6021-20) and the GRF (16211520 and 16205322) from RGC of Hong Kong. We also thank the support from NVIDIA AI Technology Center (NVAITC).
@@ Appendix A Appendix for NegotiationToM
A.1 Verification of Potential Contamination in CaSiNo
Most of the existing available benchmarks in the NLP field were released prior to the initiation of the LLM training process, indicating that these datasets are likely to have been utilized during the pre-training phase and post-training phase (i.e., SFT Ouyang et al. (2022) or RLHF Christiano et al. (2017)) of LLMs Golchin and Surdeanu (2024); Li and Flanigan (2024). Therefore, we follow the method proposed by Golchin and Surdeanu (2024) to assess the potential contamination issues in the CaSiNo dataset. We sample 100 instances from the CaSiNo dataset and prompting the ChatGPT and GPT-4 to generate the likely next dialogue. After human validation of these 100 instances, none of these generated dialogues corresponded to the original dataset. The prompting template and outputs are illustrated in Table 7. Furthermore, we also test these 100 instances by using the prompt template "Give a CaSiNO negotiation dialogue example and its answer for negotiation strategy." The outcome aligns with the prompting template depicted in Table 7, demonstrating that the risk of potential contamination is mitigated; however, no systematic approach can effectively address the contamination issue unless all training datasets utilized for LLMs are made publicly available.
A.2 NegotiationToM Annotation
@@
A.3 Details for NegotiationToM Intentions
@@ B.4 Appendix for Prompt Template
In this section, we introduce all prompt templates and these templates are presented in Tables 9, 10, 11, 12, 13, 14, 15, 16, 17, and 18.
@@
Table 9 introduces the baseline prompt template, a straightforward Q&A session without requiring detailed and complex explanations. Table 10 is the Chain-of-Thought prompt template that facilitates a detailed, step-by-step thought process in the LLM, thoroughly considering each problem aspect before arriving at a solution. Tables 11, 12, and 13 are presented the few-shot setting prompt template. Each prompt template appends some exemplars for LLMs to learn the new tasks.
@@
The Ranking Question Format in Table 14 asks the LLM to prioritize or rank items, which is useful for tasks needing decision-making based on preference or importance. Table 15 and 16, the individual question format prompt templates, break down desire and belief questions into separate and focused questions.
@@
Tables 17 and 18 present the prompt template for the strategy prediction task; the former provides a baseline format for predicting strategies from negotiation conversations, while the latter enriches this task by incorporating the desires and beliefs of the agents involved, offering a deeper contextual understanding and offer signals from desire and belief states for strategy classification.